OpenStudy (anonymous):

24a - 22 = -4(1-6a) just double checking but, there is no real answer for this, right?

jimthompson5910 (jim_thompson5910):

That is correct

jimthompson5910 (jim_thompson5910):

24a - 22 = -4(1-6a) 24a - 22 = -4+24a 24a - 22 - 24a = -4 + 24a - 24a -22 = -4 which is always false, so there are no solutions

OpenStudy (anonymous):

does it make a difference if i write no answer or no real answer or is it only no solutions

jimthompson5910 (jim_thompson5910):

no answer, no real answer, no solutions, no real solutions are 4 ways to say the same thing really

OpenStudy (anonymous):

okay good